OpenStudy (anonymous):

calculate the electric field at one corner of a square 1.7 m on the side if the other three corners are occupied by 3.05*10^-6 C charges

OpenStudy (anonymous):

You just need to figure out how far each charge is.

OpenStudy (anonymous):

\[ \large \vec{E} = k_e\sum \frac{q_i}{|\hat{r_i}|^2} \hat{r_i} \]

OpenStudy (anonymous):

For each case \(q_i = 3.05\times 10^{-6}C \) whereas \(\hat{r_1} = 1.7 m\hat{i} \), and \(\hat{r_2} = 1.7 m\hat{j} \), \(\hat{r_3} = 1.7 m\hat{i}+1.7 m\hat{j} \)